# CARLA: Self-supervised Contrastive Representation Learning for Time Series Anomaly Detection

### v3

@@ 3.2 Pretext Stage
The Pretext Stage of CARLA consists of two parts: Part One is a contrastive representation learning step that uses a ResNet architecture (which has shown effectiveness in times series classification tasks Ismail Fawaz et al. (2019)) to learn representations for time series windows, and Part Two is a post-processing step that uses the learned representations from Part One to identify semantically meaningful nearest and furthest neighbours for each learned window representation. Algorithm 2 shows the Pretext steps.
@@ 3.2.1 Part One: Contrastive Representation Learning
Our approach empowers the model to learn similar feature representations for temporally proximate windows. Since the majority of data is normal, the model captures temporal relationships of normal data through learning similar representations for normal windows that are temporally proximate. Furthermore, by introducing anomalies into the system, the model can learn a more effective decision boundary, resulting in a reduced false positive rate and enhanced precision in identifying anomalies when compared to current state-of-the-art models.
@@ 3.2.2 Part Two: Nearest and Furthest Neighbours
Part Two of our approach uses the feature representations learned in Part One to identify semantically meaningful nearest and furthest neighbours for each sample. To achieve this, we utilise $\mathcal{B}$ along with the indices of $Q$ nearest and $Q$ furthest neighbours of all samples in $\mathcal{B}$.
@@ 3.3 Self-supervised Classification Stage
To encourage class diversity and prevent overfitting, we use entropy loss on the distribution of anchor and neighbour samples across classes. Assuming the classes set is denoted as $\mathcal{C}=\{1,...,C\}$, and the probability of window $w_{i}$ being assigned to class $c$ is denoted as $\phi_{s}^{c}(w_{i})$:
@@
The goal of the loss function is to learn a representation that is highly discriminative, with nearest neighbours assigned to the same class as the anchor samples, and furthest neighbours assigned to a distinct class. By incorporating the semantically meaningful nearest and furthest neighbours, the model is able to produce more consistent and confident predictions, with the probability of a window being classified as one particular class is close to 1 or 0.
@@ 4 Experiments
Benchmark # datasets # dims Train size Test size Anomaly ratio
@@ 4.1 Benchmark Datasets
We evaluate the performance of the proposed model and make comparisons of the results across the seven most commonly used real benchmark datasets for TSAD. The datasets are summarised in Table 1, which provides key statistics for each dataset.
@@
NASA DatasetsMars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl – (Hundman et al., 2018) are real-world datasets collected from NASA spacecraft. These datasets contain anomaly information derived from reports of incident anomalies for a spacecraft monitoring system. MSL and SMAP comprises of 27 and 55 datasets respectively and each equipped with a predefined train/test split, where, unlike other datasets, their training set are unlabeled.
@@
Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset (Su et al., 2019) is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days. During this period, normal data was observed within the initial 5 days, while anomalies were sporadically injected during the subsequent 5 days. The dataset is also equipped with a predefined train/test split, where the training data is unlabeled.
@@
Secure water treatment (SWaT) [^3] https://itrust.sutd.edu.sg/testbeds/secure-water-treatment-swat/ (Mathur and Tippenhauer, 2016) is meticulously gathered within 11 days, utilising a water treatment platform that consisted of 51 sensors. During the initial 7 days, the dataset exclusively comprised normal data, accurately reflecting regular operational conditions. However, in the subsequent 4 days, a total of 41 anomalies were deliberately generated using a wide range of attack methods.
@@
Water distribution testbed (WADI) [^4] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi/ (Ahmed et al., 2017) is acquired from a scaled-down urban water distribution system that included a combined total of 123 actuators and sensors. This data covers a span of 16 days. Notably, the final two days of the dataset specifically feature anomalies. In addition, the test dataset encompasses 15 distinct anomaly segments, providing valuable examples for anomaly detection and analysis purposes.
@@
Yahoo [^5] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 dataset (Laptev et al., 2015) is a benchmark dataset that consists of hourly 367 sampled time series with anomaly points labels. The dataset includes four benchmarks, with the A1 benchmark containing ”real” production traffic data from Yahoo properties, and the other three are synthetic. For evaluation purposes, we focus on the A1 benchmark, which includes 67 univariate time series. However, some of these contain no anomalies in the test, hence we exclude them from the evaluation.
@@
Key performance indicators (KPI) [^6] https://github.com/NetManAIOps/KPI-Anomaly-Detection contains time series data of service and machine key performance indicators (KPIs) from real scenarios of Internet companies. Service KPIs measure web service performance, while machine KPIs reflect machine health. Examples of service KPIs include page response time and page views, while machine KPIs include CPU and memory utilisation. The dataset includes 29 different KPI datasets, with separate training and test datasets provided.
@@ 4.2 Benchmark Methods
In this overview, provide a description of the ten prominent TSAD models that were used for comparison with CARLA, spanning various TSAD categories. For the semi-supervised approach, the LSTM-VAE (Park et al., 2018) is highlighted, which is a point-wise anomaly detection model. In the realm of unsupervised reconstruction-based models, Donut (Xu et al., 2018) stands out for UTS, while OmniAnomaly Su et al. (2019) and AnomalyTransformer (Xu et al., 2021a) are noted for MTS, alongside TranAD (Tuli et al., 2022) which caters to both UTS and MTS. The unsupervised forecasting-based model, THOC Shen et al. (2020), is recognized for its predictive capabilities. Within the self-supervised category, MTAD-GAT (Zhao et al., 2020) emerges as a hybrid model, with TS2Vec Yue et al. (2022) and DCdetector Yang et al. (2023) offering representation-based methodologies. Additionally, recent advancements are represented by TimesNet Wu et al. (2023) and DCdetector Yang et al. (2023), showcasing the ongoing evolution in TSAD models.
Donut  [^7] https://github.com/NetManAIOps/donut Xu et al. (2018) is a method for detection anomalies in UTS. It utilises a VAE for a reconstruction-based strategy. By learning the latent representation of time series data, it accurately identifies anomalies when data reconstruction significantly deviates from the original.
LSTM-VAE [^8] https://github.com/TimyadNyda/Variational-Lstm-Autoencoder Park et al. (2018) is a point-wise model that combines LSTM networks, which capture temporal data dependencies, with VAEs for generating high-dimensional data distributions and modelling complex data patterns. This integration allows for effective anomaly detection in MTS points by identifying deviations from normal behaviour, which is crucial in applications like robot-assisted feeding for ensuring safety and efficiency.
OmniAnomaly [^9] https://github.com/smallcowbaby/OmniAnomaly Su et al. (2019) is an unsupervised model that utilises a Variational Autoencoder (VAE) to model multivariate time series data. By analyzing the likelihood of a given window’s reconstruction, the model can detect anomalies.
MTAD-GAT [^10] https://github.com/mangushev/mtad-gat Zhao et al. (2020) enhances MTS anomaly detection by leveraging graph attention networks to understand complex dependencies across different time series. It employs a self-supervised framework that uses graph attention layers to learn relationships in both temporal and feature dimensions. This approach optimizes both forecasting-based and reconstruction-based models for anomaly detection.
THOC [^11] We utilised the authors’ shared implementation, as it is not publicly available. Shen et al. (2020) utilises a multi-layer dilated recurrent neural network (RNN) alongside skip connections in order to effectively handle contextual information. THOC utilise a temporal hierarchical one-class network for anomaly detection.
AnomalyTransformer [^12] https://github.com/thuml/Anomaly-Transformer Xu et al. (2021a) introduces an approach for detecting anomalies in MTS by leveraging the Transformer architecture, known for its effectiveness in capturing long-range dependencies. It employs association discrepancy to identify anomalies, focusing on the differences in relationships among data points in a time series.
TranAD [^13] https://github.com/imperial-qore/TranAD Tuli et al. (2022) leverages deep transformer networks for efficient anomaly detection in MTS and UTS. It employs attention-based sequence encoders for quick inference, leveraging focus score-based self-conditioning and adversarial training for robust feature extraction and stability.
TS2Vec [^14] https://github.com/yuezhihan/ts2vec Yue et al. (2022) is an unsupervised contrastive model that is capable of learning multiple contextual representations of MTS and UTS semantically at various levels. TS2Vec employs contrastive learning in a hierarchical way, which provides a contextual representation. They proposed a method to be employed for time series anomaly detection.
DCdetector [^15] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al. (2023) introduces an approach for time series anomaly detection by employing dual attention mechanisms and contrastive representation learning. This method focuses on capturing intricate temporal relationships and distinguishing between normal and anomalous patterns in time series data.
TimesNet [^16] https://github.com/thuml/TimesNet Wu et al. (2023) proposes a framework for analyzing general time series data by modelling temporal 2D variations. This approach allows for capturing complex patterns and dependencies in time series. This method leverages advanced neural networks to enhance analysis accuracy and efficiency across various applications like time series anomaly detection.
Random Anomaly Score. The “Random Anomaly Score” model is designed to generate anomaly scores for time series windows based on a normal distribution ($\mathcal{N}(0,1)$) with mean 0 and standard deviation 1. This model has been specifically developed to showcase evaluation metrics and strategies in the field of TSAD.
@@ 4.3 Evaluation Setup
In our study, we evaluate ten TSAD models on benchmark datasets previously mentioned in section 4.1 using their best hyper-parameters as they stated to ensure a fair evaluation. We summarise the default hyper-parameters used in our implementation as follows: The CARLA model consists of a 3-layer ResNet architecture with three different kernel sizes [8, 5, 3] to capture temporal dependencies. The dimension of the representation is 128. We use the same hyper-parameters across all datasets to evaluate CARLA: window size = 200, number of classes = 10, number of nearest/furthest neighbours ($Q$) = 5, and coefficient of entropy = 5. For detailed information about all experiments involving the aforementioned hyper-parameter choices, please refer to the section 4.5. We run the Pretext Stage for 30 epochs and the Self-supervised Classification Stage for 100 epochs on all datasets.
It is important to note that we do not use Point Adjustment (PA) in our evaluation process. Despite its popularity, we do not use PA due to the findings in Kim et al. (2022) that its application leads to an overestimation of a TSAD model’s capability and can bias results in favour of methods that produce extreme anomaly scores. Thus, to ensure the accuracy and validity of our results, we relegate PA results to B and present here conventional F1 scores. The F1 score calculated without considering the PA is referred to as F1, while the adjusted predictions are denoted as $F1_{PA}$ in B.
Table 2: Precision (Prec), recall (Rec), F1 and AU-PR results for various models on multivariate time series datasets. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline. Due to the single time series in both the SWaT and WADI datasets, the standard deviation of AU-PR is not available.
@@
Table 3: Precision (Prec), recall (Rec), F1, and AU-PR results for various models on univariate time series datasets. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
@@ 4.4 Benchmark Comparison
The performance metrics employed to evaluate the effectiveness of all models consist of precision, recall, the traditional F1 score (F1 score without PA), FPR and the area under the precision-recall curve (AU-PR). Additionally, we computed the average ranks of the models based on their F1. For all methods, we used the precision-recall curve on the anomaly score for each time series in the benchmark datasets to find the best F1 score based on precision and recall for the target time series.
@@
Furthermore, for datasets with more than one time series, we report the average and standard deviation of AU-PR across all time series. AU-PR is advantageous in imbalanced datasets because it is less sensitive to the distribution of classes Saito and Rehmsmeier (2015). It considers the trade-off between precision and recall across all possible decision thresholds, making it more robust in scenarios where the number of instances in classes are imbalanced. By focusing on the positive class (i.e. anomalous) and its predictions, AU-PR offers a more precise evaluation of a model’s ability to identify and prioritise the minority class correctly.
@@
Tables 2 and 3 demonstrate the performance comparison between CARLA and all benchmark methods for multivariate and univariate datasets, respectively. These two tables show our model outperforms other models with higher F1 and AU-PR scores across all datasets (except SWaT for F1 and AU-PR and Yahoo-A1 for AU-PR), in which CARLA is the third best. This shows the strength of CARLA in generalising normal patterns due to its high precision. CARLA’s precision is the highest across all seven datasets. At the same time, CARLA has high enough recall, and as a result, our CARLA’s F1 is the highest across all datasets except SWaT. While the recall is higher in other benchmarks, such as OmniAnomaly, it is with the expense of very low precision (with median precision $<19.67\%$ across all datasets). This means that a good balance is achieved between precision and recall by CARLA. This is also shown in our consistently high AU-PR compared to other models and indicates that it is proficient in accurately identifying anomalous instances with higher recall while minimising false alarms with higher precision on imbalanced datasets where normal instances are predominant. Finally, CARLA achieved the lowest average rank (best rank), based on F1, in both multivariate and univariate benchmarks (see last column in Tables 2 and 3).
@@
Figure 4: Models’ performance comparison on multivariate time series datasets regarding their FPR and AU-PR. Models in the optimal quadrant (low FPR, high AU-PR) demonstrate superior anomaly detection with minimal false alarms.
Figure 5: Models’ performance comparison on univariate time series datasets regarding their FPR and AU-PR. Models in the optimal quadrant (low FPR, high AU-PR) demonstrate superior anomaly detection with minimal false alarms.
@@
We now evaluate the performance of all models in terms of FPR. We also incorporate AU-PR in this comparison, as some models can achieve a low FPR at the expense of lowering AU-PR. In the scatter plots provided in Figure 4 we show FPR on the x-axis and AU-PR on the y-axis. The ideal point on this scatter plot is in the leftmost top corner. CARLA consistently shows a lower FPR compared to other models across various multivariate time series datasets. This is particularly noteworthy in the context of anomaly detection, where the cost of false positives can be substantial, leading to wasted resources and potential overlook of true anomalies. CARLA’s position in the plots are close to leftmost top corner (in the optimal quadrant). This indicates fewer false alarms while maintaining competitive AU-PR scores, which measures the precision and recall balance of anomaly detection.
@@
For instance, on the MSL dataset, CARLA is among the models with the lowest FPR, and its AU-PR score is in the upper tier, showcasing its ability to identify true anomalies accurately. On the SMAP and SMD datasets, CARLA maintains a significantly lower FPR than most models, suggesting its robustness to varying data conditions. In the SWaT dataset, CARLA excels with the lowest FPR and the third highest AU-PR, marking it suitable for precision-critical applications. Even on WADI, while CARLA’s FPR is slightly higher, it remains in the optimal quadrant, reflecting a strong FPR and AU-PR balance vital for maintaining operational continuity and effective anomaly detection.
@@
Adding to CARLA’s impressive performance on multivariate datasets, its capabilities in univariate time series analysis, as shown in Figure 5, further highlighting its robustness. On the Yahoo-A1 dataset, while CARLA does not achieve the highest AU-PR, it still demonstrates a commendable balance of a low FPR and a high AU-PR, positioning it as the third-best model in this regard. Its place in the upper left quadrant indicates a strong ability to correctly identify anomalies with a minimal rate of false alarms. On the KPI dataset, CARLA maintains a competitive edge, presenting a low FPR that surpasses most other models. While the AU-PR is the highest, CARLA’s ability to limit false positives is a notable strength.

### v4

@@ 3.2 Pretext Stage
The Pretext Stage of CARLA consists of two parts: Part One is a contrastive representation learning step that uses a ResNet architecture (which has shown effectiveness in times series classification tasks [13]) to learn representations for time series windows, and Part Two is a post-processing step that uses the learned representations from Part One to identify semantically meaningful nearest and furthest neighbours for each learned window representation. Algorithm 2 shows the Pretext steps.
@@ 3.2.1 Part One: Contrastive Representation Learning
Our approach empowers the model to learn similar feature representations for temporally proximate windows. Since the majority of data is normal, the model captures temporal relationships of normal data through learning similar representations for normal windows that are temporally proximate. Furthermore, by introducing anomalies into the system, the model can learn a more effective decision boundary, resulting in a reduced false positive rate and enhanced precision in anomaly detection compared to current state-of-the-art models.
@@ 3.2.2 Part Two: Nearest and Furthest Neighbours
Part Two of our approach uses the feature representations learned in Part One to identify semantically meaningful nearest and furthest neighbours of each sample. To achieve this, we utilise $\mathcal{B}$ along with the indices of $Q$ nearest and $Q$ furthest neighbours of all samples in $\mathcal{B}$.
@@ 3.3 Self-supervised Classification Stage
To encourage class diversity and prevent overfitting, we apply entropy loss on the distribution of anchor and neighbour samples across classes. Assuming the classes set is denoted as $\mathcal{C}=\{1,...,C\}$, and the probability of window $w_{i}$ being assigned to class $c$ is denoted as $\phi_{s}^{c}(w_{i})$:
@@
The goal of the loss function is to learn a representation that is highly discriminative, with the nearest neighbours assigned to the same class as the anchor samples and the furthest neighbours assigned to a distinct class. By incorporating the semantically meaningful nearest and furthest neighbours, the model is able to produce more consistent and confident predictions, with the probability of a window being classified as one particular class is close to 1 or 0.
@@ 4 Experiments
Benchmark # datasets # dims Train size Test size Anomaly%
@@ 4.1 Benchmark Datasets
We evaluate the performance of the proposed model and make comparisons of the results across the seven most commonly used real benchmark datasets for TSAD. The datasets are summarised in Table 1, which provides key statistics for each dataset. All datasets, except Yahoo, have a predefined train/test split with unlabeled training data.
@@
NASA DatasetsMars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl – [12] These real-world datasets, collected from NASA spacecraft, contain anomaly information from incident reports for a spacecraft monitoring system.
@@
Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset [27] is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days. During this period, normal data was observed within the initial 5 days, while anomalies were sporadically injected during the subsequent 5 days.
@@
Secure water treatment (SWaT) [^3] https://itrust.sutd.edu.sg/testbeds/secure-water-treatment-swat/ [19] is meticulously gathered over 11 days, utilising a water treatment platform with 51 sensors. During the initial 7 days, it consisted exclusively of normal data, reflecting regular operational conditions. In the following 4 days, 41 anomalies were deliberately generated using a wide range of attacks.
@@
Water distribution testbed (WADI) [^4] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi/ [1] is acquired from a scaled-down urban water distribution system that included a combined total of 123 actuators and sensors. This data covers a span of 16 days. Notably, the final two days of the dataset specifically feature anomalies. In addition, the test dataset encompasses 15 distinct anomaly segments, providing valuable examples for anomaly detection and analysis purposes.
@@
Yahoo [^5] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 dataset [15] is a benchmark dataset that consists of hourly 367 sampled time series with anomaly points labels. The dataset includes four benchmarks, with the A1 benchmark containing ”real” production traffic data from Yahoo properties, and the other three are synthetic. For evaluation, we focus on the A1 benchmark, which includes 67 univariate time series. We use half of each dataset for training and the remaining for test, excluding those without anomalies in the test set.
@@
Key performance indicators (KPI) [^6] https://github.com/NetManAIOps/KPI-Anomaly-Detection contains data of service and machine key performance indicators (KPIs) from real scenarios of Internet companies. Service KPIs measure web service performance, such as page response time and page views. Machine KPIs reflect machine health, including CPU and memory utilisation.
@@ 4.2 Benchmark Methods
We provide a description of the ten prominent TSAD models that were used for comparison with CARLA, spanning various TSAD categories. For the semi-supervised approach, the LSTM-VAE [23] is highlighted, which is a point-wise anomaly detection model. In the realm of unsupervised reconstruction-based models, Donut [31] stands out for UTS, while OmniAnomaly [27] and AnomalyTransformer [34] are noted for MTS, alongside TranAD [29] which caters to both UTS and MTS. The unsupervised forecasting-based model, THOC [26] and TimesNet [30], are recognised for their predictive capabilities. Within the self-supervised category, MTAD-GAT [40] emerges as a hybrid model, with TS2Vec [38] and DCdetector [36] offering representation-based methodologies. Additionally, recent advancements are represented by TimesNet [30] and DCdetector [36], showcasing the ongoing evolution in TSAD models. The “Random Anomaly Score” model is designed to generate anomaly scores based on a normal distribution ($\mathcal{N}(0,1)$) with mean 0 and standard deviation 1. This model is specifically developed to illustrate evaluation metrics and strategies in the field of TSAD.
@@ 4.3 Evaluation Setup
In our study, we evaluate ten TSAD models on benchmark datasets previously mentioned in section 4.1 using their best hyper-parameters, as they stated, to ensure a fair evaluation. We summarise the default hyper-parameters used in our implementation as follows: The CARLA model consists of a 3-layer ResNet architecture with three different kernel sizes [8, 5, 3] to capture temporal dependencies. The dimension of the representation is 128. We use the same hyper-parameters across all datasets to evaluate CARLA: window size = 200, number of classes = 10, number of nearest/furthest neighbours ($Q$) = 5, and coefficient of entropy = 5. For detailed information about all experiments involving the aforementioned hyper-parameter choices, please refer to the section 4.5. We run the Pretext Stage for 30 epochs and the Self-supervised Classification Stage for 100 epochs on all datasets.
It is important to note that we do not use Point Adjustment (PA) in our evaluation process. Despite its popularity, [14] found that applying PA leads to an overestimation of TSAD models’ capability and can bias results towards methods that produce extreme anomaly scores. To ensure accuracy, we present conventional F1 scores and relegate PA results to B. The F1 score without PA is referred to as F1.
Table 2: Precision (Prec), recall (Rec), F1 and AU-PR results for various models on multivariate time series datasets. The best results are in bold, and the second-best is indicated by underline. Due to the single time series in both the SWaT and WADI datasets, the standard deviation of AU-PR is not available.
@@
Table 3: Precision (Prec), recall (Rec), F1, and AU-PR results for various models on univariate time series datasets. The best results are in bold, and the second-best is indicated by underline.
@@ 4.4 Benchmark Comparison
The performance metrics employed to evaluate the effectiveness of all models consist of precision, recall, the traditional F1 score (F1 score without PA), FPR and the area under the precision-recall curve (AU-PR). Additionally, we computed the average ranks of the models based on their F1. For all methods, we used the precision-recall curve on the anomaly score for each time series in the datasets to find the best F1 score based on precision and recall for the target time series.
@@
Furthermore, for datasets with more than one time series, we report the average and standard deviation of AU-PR across all time series. AU-PR is advantageous in imbalanced datasets because it is less sensitive to the distribution of classes [24]. It considers the trade-off between precision and recall across all possible decision thresholds, making it more robust in scenarios where the number of instances in classes is imbalanced. By focusing on the positive class (i.e. anomalous) and its predictions, AU-PR offers a more precise evaluation of a model’s ability to identify and prioritise the minority class correctly.
Figure 4: Critical difference diagrams for multivariate (4(a)) and univariate (4(b)) time series benchmark datasets.
@@
Tables 2 and 3 demonstrate the performance comparison between CARLA and all benchmark methods for multivariate and univariate datasets, respectively. These two tables show our model outperforms other models with higher F1 and AU-PR scores across all datasets (except SWaT for F1 and AU-PR and Yahoo-A1 for AU-PR), in which CARLA is the third best. This shows the strength of CARLA in generalising normal patterns due to its high precision. CARLA’s precision is the highest across all seven datasets. At the same time, CARLA has high enough recall, and as a result, our CARLA’s F1 is the highest across all datasets except SWaT. While the recall is higher in other benchmarks, such as OmniAnomaly, it is with the expense of very low precision (with median precision $<19.67\%$ across all datasets). This means that a good balance between precision and recall is achieved by CARLA. This is also shown in our consistently high AU-PR compared to other models and indicates that it is proficient in accurately identifying anomalous instances with higher recall while minimising false alarms with higher precision on imbalanced datasets where normal instances are predominant. Finally, CARLA achieved the lowest average rank (best rank), based on F1, in both multivariate and univariate benchmarks (see last column in Tables 2 and 3).
Figure 4 shows the average ranking for each model in a critical difference diagram [6]. Models within the same clique (black bar) are not statistically significant. Statistically, the Wilcoxon signed-rank test with Holm correction is used as a post hoc test to Friedman’s test. The diagrams cover 208 univariate and multivariate datasets (detailed in Table 1). CARLA consistently ranks as the best-performing model, scoring 1.00 in univariate datasets (Figure 4(b)) and 1.60 in multivariate datasets (Figure 4(a)), highlighting its superiority. Moreover, CARLA is significantly different from other models.
@@
Figure 5: Models’ performance comparison on MTS & UTS datasets regarding their FPR and AU-PR. Models in the optimal quadrant (low FPR, high AU-PR) have superior anomaly detection with minimal false alarms.
@@
We now evaluate the performance of all models in terms of FPR. We also incorporate AU-PR in this comparison, as some models can achieve a low FPR at the expense of lowering AU-PR. In the scatter plots provided in Figure 5, we show FPR on the x-axis and AU-PR on the y-axis. The ideal point on this scatter plot is in the leftmost top corner. CARLA consistently shows a lower FPR compared to other models across various multivariate time series datasets. This is particularly noteworthy in the context of anomaly detection, where the cost of false positives can be substantial, leading to wasted resources and potential overlook of true anomalies. CARLA’s position in the plots is close to the leftmost top corner (in the optimal quadrant). This indicates fewer false alarms while maintaining competitive AU-PR scores, which measure the precision and recall balance of anomaly detection.
@@
For instance, on the MSL dataset, CARLA is among the models with the lowest FPR, and its AU-PR score is in the upper tier, showcasing its ability to identify true anomalies accurately. On the SMAP and SMD datasets, CARLA maintains a significantly lower FPR than most models, suggesting its robustness to varying data conditions. In the SWaT dataset, CARLA excels with the lowest FPR and the third highest AU-PR, marking it suitable for precision-critical applications. Even on WADI, while CARLA’s FPR is slightly higher, it remains in the optimal quadrant, reflecting a strong FPR and AU-PR balance vital for operational continuity and effective anomaly detection.
@@
Adding to CARLA’s impressive performance on multivariate datasets, its capabilities in univariate time series analysis, as shown in Figure 5, further highlighting its robustness. On the Yahoo-A1 dataset, while CARLA does not achieve the highest AU-PR, it still demonstrates a commendable balance of a low FPR and a high AU-PR, positioning it as the third-best model in this regard. Its place in the upper left quadrant indicates a strong ability to correctly identify anomalies with a minimal rate of false alarms. On the KPI dataset, CARLA maintains a competitive edge with a low FPR that surpasses most other models. While the AU-PR is the highest, CARLA’s ability to limit false positives is a notable strength.